ASSIGNMENT OF THE ν_{9} FUNDAMENTAL OF THE CIS CONFORMER OF METHYL NITRITE IN THE HIGH RESOLUTION FTIR SPECTRUM

\underline{LISA M. GOSS} , Department of Chemistry, Idaho State University, Campus Box 8023, Pocatello ID, 83209; THOMAS A. BLAKE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99352.

The infrared spectrum of gas phase methyl nitrite (CH_{3}ONO) has been recorded at high resolution between 600 and 4000 cm^{−1}. Methyl nitrite contains a mixture of cis and trans rotational isomers which are interconverted by the N=O torsion. As in the ν_{8} vibration of the cis form, the ν_{9} vibration is not too congested, consistent with the high barrier to rotation of the methyl group in the cis form. The ν_{9}, with an origin at 627.9 cm ^{−1}, has been assigned by previous workers to the ONO bend of the cis conformer and several hot bands are also present in this region. In this study, a spectrum taken at 0.0015 cm ^{−1} resolution on the PNNL IFS120/HR in a 20 cm cell was used to assign more than 500 A-type transitions of the cis ν_{9} between 600 and 650 cm ^{−1} and a set of excited state constants was determined.